# Recursive Deep Inverse Reinforcement Learning

### v4

@@ 5 Recursive Deep Inverse Reinforcement Learning
In the previous section, we derived the upper bound of the negative log-likelihood cost described in Equation (2). In this section, we seek to minimize the upper bound (10) recursively. To do so, we re-write the EKF optimization problem using the upper bound derived in (10) and a regularization term. Given an expert trajectory $\mathcal{\bm{D}}_{\mathrm{demo}}\triangleq\{\tau^{(0)},\dots,\tau^{(N-1)}\}$ we seek to determine an optimal solution $\theta^{*}(t_{i})$ starting from initial condition $\theta(t_{0})$ by solving the following mathematical optimization function:
$\mathcal{L}_{N}(\Theta_{N})$ $=\mathcal{L}_{UB}+\frac{1}{2}\sum_{i=1}^{N}\lVert\theta(t_{i})-\theta(t_{i-1})\rVert^{2}_{Q_{\theta}^{-1}}$ (11)
@@
where the second term in the right-hand side of (11) is a regularization term typical to Bayesian filtering algorithms Imbiriba et al. (2022); Ghanem et al. (2025). In a similar fashion to Kalman filtering optimization process described in (Humpherys et al., 2012; Ghanem et al., 2023), we seek to determine optimal solution $\Theta_{N}^{*}=\{\theta^{*}(t_{0}),\dots,\theta^{*}(t_{N})\}$ using the second-order Newton method sequentially, which recursively finds $\Theta_{N}^{*}$ given $\Theta_{N-1}^{*}$. Noticing that problem (11) can be broken into predictive and update problems, we can derive its recursive solution, which is detailed in Section A.4 of the supplemental document, and leads to the result in Theorem 5.1.
@@
using Lemma B.3 in (Humpherys et al., 2012), the lower block $P_{{\theta}_{i}}$ of $\left(\nabla^{2}\mathcal{L}_{i}(\hat{\Theta}_{i|i-1})\right)^{-1}$ is calculated as in (13) ∎
As a consequence of Theorem (5.1), $\hat{\theta}(t_{i})$ is computed according to (12) using $\hat{\theta}(t_{i-1})$. The entire training procedure is detailed in Algorithm 1, where a detailed description of Algorithm 1 is described in Section A.3 of the supplemental document.
@@ 6 Experiments
In our experiments, we seek to answer two questions: $(i)$ Is the Upper Bound (UB) derived in (10) tight enough and useful? $(ii)$ Does RDIRL outperforms state-of-the-art methods and traditional optimization approaches of (10) in complex continuous control and adversarial tasks?
To answer $(i)$, we compare the UB in (10) using Adam optimizer in challenging continuous control scenarios from openAI gym(Brockman, 2016) and MuJoCo(Todorov et al., 2012). For RDIRL, the reward function of the benchamrked environments is learned online according to Algorithm 1 with $q(\tau)$ being an MPPI control policy.
We compare UB with state-of-the-art IRL and imitation learning algorithms such as GAIL (Ho and Ermon, 2016), Guided Cost Learning (GCL) (Finn et al., 2016b) the Adversarial Inverse Reinforcement Learning (AIRL) algorithm proposed by (Fu et al., 2017), SQIL (Reddy et al., 2019) and our proposed approach RDIRL. We find that UB matches it’s original cost function slightly outperforms GCL in CartPole, MountainCar and Hopper while it slightly underperfoms against it in the HalfCheetah-v4 environment. We show that the upper bound derived in (10) is tight enough and useful as a result.
To answer $(ii)$, we evaluate the RDIRL in continuous control scenarios from openAI and MuJoCo and a in an adversarial cognitive radar scenario (Potter et al., 2024; Haykin, 2006).
We compare RDIRL against against state of the art IRL and imitation learning algorithms such as GAIL, GCL, AIRL, SQIL (Reddy et al., 2019) in addition to our upper bound (10) UB. We find that RDIRL outperforms all of the benchmarked methods in recovering reward functions. We show that reward function functions learned with RDIRL produce optimal or near-optimal behavior much faster than competing methods. Since competing methods do not consider online adaptation, they need more expert observations and environment steps to converge.
Table 1: Comparison of mean reward values for different Gym environments and methods.
@@
SQIL (Reddy et al., 2019) $142.12\pm 13.24$ $-6.68\pm 0.34$ $-76.83\pm 43.07$ $\bm{257.15\pm 41.05}$
GAIL (Ho and Ermon, 2016) $140.22\pm 8.75$ $1725.59\pm 1479.63$ $-26.09\pm 58.04$ $194.07\pm 21.72$
GCL (Finn et al., 2016b) $138.08\pm 13.83$ $1806.03\pm 1441.38$ $-11.34\pm 64.05$ $187.37\pm 20.23$
AIRL (Fu et al., 2017) $142.96\pm 10.39$ $1698.28\pm 1489.73$ $-27.13\pm 55.62$ $192.68\pm 22.93$
UB (Upper Bound) $140.78\pm 14.06$ $1846.34\pm 1444.68$ $-16.12\pm 56.18$ $194.27\pm 20.35$
RDIRL (ours) $\bm{149.02\pm 2.09}$ $\bm{4862.09\pm 2408.55}$ $\bm{47.16\pm 66.72}$ $220.79\pm 28.05$
@@ 6.1 Continuous control
To assess the performance of our proposed approach RDIRL and the upper bound tightness, we conduct inverse reinforcement learning (IRL) experiments on the CartPole and Mountain Car environments from OpenAI Gym (Brockman, 2016) and HalfCheetah-v4, Hopper, and Walker2d from MuJoCo(Todorov et al., 2012), all solved using model-free reinforcement learning. Each task has a predefined true reward function provided by OpenAI Gym.
@@
We first generate expert demonstrations for these tasks by training a PPO reinforcement learning agent (Schulman et al., 2017) to maximize the true reward function. Each expert demonstration consists of a state trajectory of size $N$ steps specified in Table 3 in A.1 for each task, which is then used as the sole expert trajectory for each IRL algorithm. Note that we do not use expert control sequences trajectory since we do not have access to the expert’s control policy.
@@
Next, we execute RDIRL to learn the reward function and train competing IRL algorithms using the expert trajectory over multiple episodes, where each episode consists of an expert trajectory. This process is repeated for 12 Monte Carlo runs with different seeds. In all experiments, we use MPPI as the internal control policy $q(\tau)$ to maximize the learned reward function, $-c_{\theta}$. A detailed experiment description and parameter values of MPPI and IRL algorithms is described in supplemental document A.1
@@
We compare RDIRL ’s performance against GAIL (Ho and Ermon, 2016), GCL (Finn et al., 2016b), and AIRL(Fu et al., 2017),SQIL (Reddy et al., 2019) and upper bound UB. For each algorithm, we plot the cumulative true reward values of trajectories $\tau^{\mathrm{samp}}$ sampled from the inner control policy $q(\tau)$ in Figure 1. In the case of RDIRL, $\tau^{\mathrm{samp}}$ used to calculate the reward function in Figure 1 are generated online during training according to Algorithm 1. For the rest of the methods, $\tau^{\mathrm{samp}}$ are generated offline after each offline training episode is completed.
@@
All algorithms use the same neural network architecture to parameterize the reward function. Networks are randomly initialized at the start of each experiment, and all experiments are run on Nvidia-H200 GPU Cluster with 1 GPU per job(seed).
Our proposed method, RDIRL, successfully learns reward functions of the benchmarked environments and significantly outperforms the benchmark methods. In the CartPole environment, RDIRL converges within the first episode and matches the expert after the first episode, successfully recovering the expert’s true reward function. In the Mountain Car environment, it accurately learns the expert’s reward function within a short number of time steps. Similarly, RDIRL outperform competing methods in HalfCheetah environment in terms of convergence speed and reward function quality. In contrast, the other IRL methods require significantly more episodes to converge—or fail to converge altogether. We illustrate the learning curves in Figure 1 and Figure 2 with their standard deviations, demonstrating RDIRL ’s superior efficiency and performance. Note that for Walker2d, Figure 2, the learned reward values are normalized with respect to the expert reward. Moreover, similar to tasks performed in (Reddy et al., 2019) the best reward is the one that are closer to the expert. Further experimental results with detailed descriptions are provided in the supplemental document A.5.
@@
For each task, we start the expert from the same initial condition while the trained agent starts at random initial condition at each epoch for all algorithm. Moreover, Table 1 provides a comparison of mean reward values for different Gym environments and methods, highlighting the quantitative advantage of RDIRL across both discrete and continuous control tasks. Notably, RDIRL achieves the highest average rewards in most tested environments. This consistent outperformance suggests that the recursive structure and adaptive uncertainty-aware updates of RDIRL contribute to more sample-efficient and stable learning dynamics compared to traditional IRL approaches. Due to the adaptive nature of our approach, it does not require a fixed learning rate, as the matrix $P_{\theta}$ is updated at every time step and effectively acts as an adaptive learning rate.
@@ 6.2 Cognitive radar
To evaluate whether our method can learn cost functions of adversarial agents, we perform inverse reinforcement learning experiments on a cognitive radar task. The task involves a radar chasing a moving target in 3D space. The target kinematic model follows constant velocity motion (Baisa, 2020) and the radar follows a second order unicycle model (Potter et al., 2024), where the target is moving linearly in space while the radar maximizes its Fisher Information Matrix (FIM) (Potter et al., 2024) to keep track of the target. Both the radar and the target live in the same 3D $x,y,z$ Cartesian plane. The goal of the target is to learn the radar’s FIM from what it can observe from radar’s states, which is in our case radar’s position in 3D $x,y,z$ Cartesian coordinates. To achieve this goal using RDIRL,we execute Algorithm 1 where the radar’s cost function is learned online. The radar’s (expert) policy inside Algorithm 1 is an MPPI that maximizes radar’s FIM. The inner control policy $q(\tau)$ is an MPPI that maximizes the learned reward function, $-c_{\theta}$. The environmnent and IRL method’s parameters are described in Table (3) inside the supplemental document.
@@
Furthermore, we compare RDIRL against Generative Adversarial Imitation Learning (GAIL), AIRL and GCL. To implement these methods, we generate expert trajectories for multiple episodes, where the expert policy is an MPPI that maximizes the radar’s FIM. The inner control policy $q(\tau)$ in all of these baselines is an MPPI, with paramters specified in table 3. We repeat this process for 5 Monte Carlo runs using different seeds.
@@ 7 Conclusions
We presented RDIRL within the IRL framework that generalizes recent advances in maximum entropy deep IRL to online settings. We first established an equivalent upper bound loss function in Equation (10) to the negative log likelihood in Equation (2) to make recursive training of maximum entropy IRL feasible with deep learning cost function representation. Second, we leveraged sequential second-order Newton optimization to derive an online IRL algorithm by minimizing the upper bound of (2) recursively and therefore established key theoretical properties of maximum entropy online deep IRL
@@
Acknowledgments.
Owen L. Howell is grateful to the National Science Foundation Graduate Research Fellowship Program (NSF-GRFP) for financial support.
Appendix A Experiment details
@@
A.1 Training
In all our experiments, we use MPPI [Williams et al., 2016] as inner policy $q(\tau)$ in our baseline methods. MPPI is a probabilistic model predictive control policy that estimates an optimal action distribution that minimizes an agent’s objective cost function. To do so, MPPI samples a number of trajectories and weighs these trajectories depending on how well they minimize the cost function, then updates the mean of its action distribution $q(\tau)$ accordingly. Since MPPI in an online policy , i.e it updates itself every time step, it makes it a natural choice of inner policy for online IRL problems, as we noticed in our preliminary experiments that it is much more stable and has faster convergence that traditional RL methods when implemented inside RDIRL.
The implementation of the baselines (GCL, AIRL,SQIL and GAIL) are adapted from available public repository[HumanCompatibleAI, 2021].In addition, we add UB as an additional baseline, where UB is the baseline that uses the upper bound loss function 10.Furthermore, we adapt all the baselines to use MPPI as inner policy alongside our proposed approach. Since the inner policy is not SAC anymore like it was in the original baselines repositories, we tune the parameters of all the adapted baselines using grid search to produce best possible performance. The resulting parameters were used directly in RDIRL and UB. We list the hyper-parameters of all the baselines used in different environments in Table 3. These hyper-parameters were selected via grid search.
@@ A.1 Training
Cartpole-v1 $41e-$ 4 150 15 150 1e-3 50 2000
MountainCar-v0 $41e-$ 4 200 15 200 1e-2 85 3500
HalfCheetah-v4 $41e-$ 4 200 15 200 1e-2 50 500
Walker2d $41e-$ 4 200 15 200 1e-2 50 500
Hopper $41e-$ 4 200 15 200 1e-2 50 500
Cognitive Radar $41e-$ 4 200 10 200 1e-2 10 25
@@
PPO [Schulman et al., 2017] is used as the base MaxEnt RL algorithm for the expert policy. Adam is used as the optimizer.
@@
A.2 Reward Function and Discriminator Network Architectures

### v5

@@ 5 Recursive Deep Inverse Reinforcement Learning
In the previous section, we derived the upper bound of the negative log-likelihood cost described in equation 2 and showed it’s equivalent to moment matching (Swamy et al., 2021). In this section, we seek to minimize the moment matching loss of equation 10 recursively. To do so, we re-write the EKF optimization problem using the loss function derived in equation 10 and a regularization term. Given an expert trajectory $\mathcal{\bm{D}}_{\mathrm{demo}}\triangleq\{\tau^{(0)},\dots,\tau^{(N-1)}\}$ we seek to determine an optimal solution $\theta^{*}(t_{i})$ starting from initial condition $\theta(t_{0})$ by solving the following mathematical optimization function:
$\mathcal{L}_{N}(\Theta_{N})$ $=\mathcal{L}_{UB-MM}+\frac{1}{2}\sum_{i=1}^{N}\lVert\theta(t_{i})-\theta(t_{i-1})\rVert^{2}_{Q_{\theta}^{-1}}$ (11)
@@
where the second term in the right-hand side of equation 11 is a regularization term typical to Bayesian filtering algorithms Imbiriba et al. (2022); Ghanem et al. (2025). In a similar fashion to Kalman filtering optimization process described in (Humpherys et al., 2012; Ghanem et al., 2023), we seek to determine optimal solution $\Theta_{N}^{*}=\{\theta^{*}(t_{0}),\dots,\theta^{*}(t_{N})\}$ using the second-order Newton method sequentially, which recursively finds $\Theta_{N}^{*}$ given $\Theta_{N-1}^{*}$. Noticing that problem equation 11 can be broken into predictive and update problems, we can derive its recursive solution, which is detailed in Section B.4 of the Appendix, and leads to the result in Theorem 5.1.
@@
using Lemma B.3 in (Humpherys et al., 2012), the lower block $P_{{\theta}_{i}}$ of $\left(\nabla^{2}\mathcal{L}_{i}(\hat{\Theta}_{i|i-1})\right)^{-1}$ is calculated as in equation 13
As a consequence of Theorem (5.1), $\hat{\theta}(t_{i})$ is computed according to equation 12 using $\hat{\theta}(t_{i-1})$. The entire training procedure is detailed in Algorithm 1, and a detailed description of Algorithm 1 is described in Section B.3 of the Appendix.
@@ 6 Experiments
We evaluate the proposed RDIRL algorithm in continuous control benchmarks from OpenAI Gym (Brockman, 2016) and MuJoCo (Todorov et al., 2012), as well as in an adversarial cognitive radar scenario (Potter et al., 2024; Haykin, 2006). We compare its performance against state-of-the-art inverse reinforcement learning and imitation learning methods, including Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016), Guided Cost Learning (GCL) (Finn et al., 2016b), Adversarial Inverse Reinforcement Learning (AIRL) (Fu et al., 2017), SQIL (Reddy et al., 2020), and Maximum Likelihood Inverse Reinforcement Learning (ML-IRL) (Zeng et al., 2022), a moment-matching variant of IRL. Experiments are conducted in two regimes: batch mode (section 6), where competing methods are trained in their standard setting with full trajectory batches, and streaming mode, where updates occur sample by sample (Appendix B.5).
Unlike reinforcement learning methods such as SAC(Haarnoja et al., 2018) or PPO(Schulman et al., 2017), which require large trajectory batches to converge and thus fail in streaming or real-time settings, our approach leverages MPPI (Williams et al., 2016) as the inner control policy. Since MPPI updates its actions at every time step, it is naturally suited for online IRL. In preliminary experiments, MPPI also provided stable performance and fast convergence unlike traditional RL policies when integrated into the RDIRL framework. Furthermore, preliminary experiments showed that competing IRL methods paired with their original RL inner policies failed to converge in streaming mode too. For consistency and fairness, we therefore adapt all competing methods to use MPPI as the inner policy in both batch and streaming comparisons.
Our results show that RDIRL consistently outperforms all benchmarked methods in recovering reward functions. Policies trained with rewards learned by RDIRL achieve optimal or near-optimal behavior significantly faster than competing approaches. Crucially, unlike existing methods which require large batches of expert trajectories and environment rollouts to converge, RDIRL leverages online adaptation. This enables efficient learning from streaming demonstrations, making it particularly well-suited for adversarial and time-limited scenarios.
Table 1: Comparison of normalized averaged reward values across all episodes for different Gym environments and methods.
@@
SQIL (Reddy et al., 2020) $947\pm 0.088 0.0$$001\pm 4.79$ $1.56\pm 0.89$ $0.799\pm 0.15$
GAIL (Ho & Ermon, 2016) $0.934\pm 0.058$ $0.236\pm 0.203$ $0.521\pm 1.15$ $0.714\pm 0.08$
GCL (Finn et al., 2016b) $0.92\pm 0.09$ $0.247\pm 0.19$ $0.226\pm 1.27$ $0.69\pm 0.075$
AIRL (Fu et al., 2017) $0.953\pm 0.069$ $0.233\pm 0.204$ $0.54\pm 1.11$ $0.709\pm 0.084$
ML-IRL(Zeng et al., 2022) $0.938\pm 0.093$ $0.253\pm 0.19$ $0.32\pm 1.12$ $0.648\pm 0.06$
RDIRL (ours) $\bm{0.993\pm 0.013}$ $\bm{0.68\pm 0.32}$ $\bm{0.496\pm 0.59}$ $\bm{0.803\pm 0.11}$
@@ 6.1 Continuous control
To assess the performance of our proposed approach RDIRL, we conduct inverse reinforcement learning (IRL) experiments on the CartPole and Mountain Car environments from OpenAI Gym (Brockman, 2016) and HalfCheetah-v4, Hopper, and Walker2d from MuJoCo(Todorov et al., 2012), all solved using model-free reinforcement learning. Each task has a predefined true reward function provided by OpenAI Gym.
@@
We first generate expert demonstrations for these tasks by training a PPO reinforcement learning agent (Schulman et al., 2017) to maximize the true reward function. Each expert demonstration consists of a state trajectory of size $N$ steps specified in Table 3 in B.1 for each task, which is then used as the sole expert trajectory for each IRL algorithm. Note that we do not use expert control sequences trajectory since we do not have access to the expert’s control policy.
@@
Next, we execute RDIRL to learn the reward function and train competing IRL algorithms using the expert trajectory over multiple episodes in batch mode, where each episode consists of an expert trajectory. This process is repeated for 12 Monte Carlo runs with different seeds. In all experiments, we use MPPI as the internal control policy $q(\tau)$ to maximize the learned reward function, $c_{\theta}$. A detailed experiment description and parameter values of MPPI and IRL algorithms is described in Appendix B.1
@@
We plot the mean of the normalized cumulative reward values across all episodes of trajectories $\tau^{\mathrm{samp}}$ sampled from the inner control policy $q(\tau)$ in Figure 1.the averaged reward values are normalized with respect to the expert reward. In the case of RDIRL, $\tau^{\mathrm{samp}}$ used to calculate the reward function in Figure 1 are generated online during training according to Algorithm 1. For the rest of the methods, $\tau^{\mathrm{samp}}$ are generated offline after each offline training episode is completed.
@@
All methods use the same neural network architecture to parameterize the reward function. Networks are randomly initialized at the start of each experiment, and all experiments are run on Nvidia-H200 GPU Cluster with 1 GPU per job(seed).
Our proposed method, RDIRL, successfully learns reward functions across all benchmark environments and consistently outperforms competing methods. In CartPole and MountainCar, it quickly recovers the expert reward even converging in one episode in CartPole, while in HalfCheetah and Hopper it achieves faster convergence and higher reward quality than baselines, many of which require far more episodes to converge or fail to converge. Learning curves in Figures 1 and 2 illustrate these improvements, with Walker2d results consistent with Figure 3 in (Reddy et al., 2020), where rewards closer to the expert indicate better performance. Furthermore, experimental results in streaming settings with detailed descriptions are provided in Appendix B.5.
@@
Table 1 further shows that RDIRL achieves the highest normalized rewards in most tasks. This consistent outperformance stems from its recursive structure and adaptive uncertainty-aware updates, which improve sample efficiency and stability. Unlike traditional IRL, our method requires no fixed learning rate, as $P_{\theta}$ is updated at each step and acts as an adaptive rate.
@@ 6.2 Cognitive radar
To evaluate whether our method can learn cost functions of adversarial agents, we perform inverse reinforcement learning experiments on a cognitive radar task. The task involves a radar chasing a moving target in 3D space. The target kinematic model follows constant velocity motion (Baisa, 2020) and the radar follows a second order unicycle model (Potter et al., 2024), where the target is moving linearly in space while the radar maximizes its Fisher Information Matrix (FIM) (Potter et al., 2024) to keep track of the target. Both the radar and the target live in the same 3D $x,y,z$ Cartesian plane. The goal of the target is to learn the radar’s FIM from what it can observe from radar’s states, which is in our case radar’s position in 3D $x,y,z$ Cartesian coordinates. To achieve this goal using RDIRL,we execute Algorithm 1 where the radar’s cost function is learned online. The radar’s (expert) policy inside Algorithm 1 is an MPPI that maximizes radar’s FIM. The inner control policy $q(\tau)$ is an MPPI that maximizes the learned reward function, $c_{\theta}$. The environmnent and IRL method’s parameters are described in Table equation 3 inside the Appendix.
@@
Furthermore, we compare RDIRL against GAIL, AIRL and GCL. To implement these methods, we generate expert trajectories for multiple episodes, where the expert policy is an MPPI that maximizes the radar’s FIM. The inner control policy $q(\tau)$ in all of these baselines is an MPPI, with parameters specified in table 3. We repeat this process for 5 Monte Carlo runs using different seeds.
@@ 7 Conclusions
We presented RDIRL within the IRL framework that generalizes recent advances in maximum entropy deep IRL to online settings. We first established the equivalence between upper bound loss function in equation 10 of the negative log likelihood in equation 2 to moment matching loss of (Swamy et al., 2021). Second, we leveraged sequential second-order Newton optimization to derive an online IRL algorithm by minimizing the moment matching loss function of equation 10 recursively and therefore established key theoretical properties of maximum entropy online deep IRL
@@
Appendix A Appendix
Appendix B Experiment details
@@
B.1 Training
In all our experiments, we use MPPI (Williams et al., 2016) as inner policy $q(\tau)$ in our baseline methods. MPPI is a probabilistic model predictive control policy that estimates an optimal action distribution that minimizes an agent’s objective cost function. To do so, MPPI samples a number of trajectories and weighs these trajectories depending on how well they minimize the cost function, then updates the mean of its action distribution $q(\tau)$ accordingly. Since MPPI in an online policy , i.e it updates itself every time step, it makes it a natural choice of inner policy for online IRL problems, as we noticed in our preliminary experiments that it is much more stable and has faster convergence that traditional RL methods when implemented inside RDIRL.
The implementation of the baselines (GCL, AIRL,SQIL,ML-IRL and GAIL) are adapted from available public repository(HumanCompatibleAI, 2021).Furthermore, we adapt all the baselines to use MPPI as inner policy alongside our proposed approach. Since the inner policy is not SAC anymore like it was in the original baselines repositories, we tune the parameters of all the adapted baselines using grid search to produce best possible performance. The resulting parameters were used directly in RDIRL. We list the hyper-parameters of all the baselines used in different environments in Table 3. These hyper-parameters were selected via grid search.
@@ B.1 Training
Cartpole-v1 1e41e-4 150 15 150 1e-3 50 2000
MountainCar-v0 1e41e-4 200 15 200 1e-2 85 3500
HalfCheetah-v4 1e41e-4 200 15 200 1e-2 50 500
Walker2d 1e41e-4 200 15 200 1e-2 50 500
Hopper 1e41e-4 200 15 200 1e-2 50 500
Cognitive Radar 1e41e-4 200 10 200 1e-2 10 25
@@
PPO (Schulman et al., 2017) is used as the base MaxEnt RL algorithm for the expert policy. Adam is used as the optimizer.
@@
B.2 Reward Function and Discriminator Network Architectures